# Bott-integrability of overtwisted contact structures

### v1

@@ 4.1.2. Introducing Lutz components and critical orbits
This modified contact form still has the same Bott integral (a monotone function of $s$), and the homotopy class of the contact structure as a tangent $2$-plane field is unchanged by the full Lutz twist; see [3, Lemma 4.5.3].
In a second step, we modify the Bott integral by a finite number of ‘reverse Giroux eliminations’ in the $(s,\varphi)$- or the $(s,\theta)$-plane transverse to the Reeb flow in the direction of ±$\pm\partial_{\theta}$ or $\partial_{\varphi}$, respectively; see [4, Section 2.3] for the details. In this way, we can introduce any finite number of pairs of elliptic and hyperbolic critical orbits into the Reeb flow, lying in the isotopy class of ±$\pm S^{1}_{\theta}$ (the fibre class) or $S^{1}_{\varphi}$ (the class represented by $\gamma$).
@@ 4.2.2. Fibre connected sum
The general construction of fibre connected sums of contact manifolds along two diffeomorphic contact submanifolds with opposite normal bundles has been described in [3, Section 7.4].
In our situation, an ad hoc construction is much simpler and yields the Bott-integrability to boot. Indeed, the fibre connected sum along two elliptic critical Reeb orbits can be defined by removing standard neighbourhoods as in [4, Theorem 2.4] around the two orbits, and then identifying the toral boundaries of the complements by a diffeomorphism that sends the fibre circle on one side to the fibre circle on the other, and glues the meridional circles with a flip of the orientation. Notice that the outer normal on one complement glues with the inner normal on the other, so the gluing of the torus boundary should indeed be done with an orientation-reversing diffeomorphism.
Since the contact form near the boundary of a tubular neighbourhood of an elliptic orbit is of Lutz type, the Bott-integrable contact form on the fibre connected sum can now be constructed by an application of the sewing lemma [4, Lemma 3.4].
@@ 5.1. Graph manifolds and the JSJ decomposition
The summands $M_{1},\ldots,M_{n}$ in the prime decomposition of a graph manifold $M$ are also graph manifolds; see Theorems 6.3 and 7.1 in [16]. Conversely, the connected sum of graph manifolds is a graph manifold; see the remark on page 91 of [16] or [4, Lemma 3.1].
@@
If $M$ is an irreducible manifold, it can be cut along a disjoint collection of incompressible tori into pieces that are either atoroidal or Seifert fibred. Rather amazingly, the minimal (with respect to inclusion) collection of such tori is unique up to isotopy; see [8, Theorem 1.9]. This unique decomposition is known as the JSJ decomposition of $M$, named after Jaco–Shalen [10] and Johannson [12].
It turns out that amongst irreducible $3$-manifolds, graph manifolds are characterised as those whose JSJ decomposition only contains Seifert fibred pieces. One direction of this statement is obvious; the other is the content of [16, Theorem 6.3]. In fact, in the contemporary literature this characterisation is often taken as the definition of a graph manifold.
@@ 5.2.1. Graph manifolds prime to $S^{1}\times S^{2}$
As a crucial preliminary step for this theorem, Yano establishes the particular case that every element in the first homology group of a Seifert manifold can be represented by a graph link [18, Proposition 2.2]. Our argument in Section 4 gives an alternative proof of that part.
@@ 5.2.2. Graph manifolds containing $(S^{1}\times S^{2})$-summands
Here is Yano’s homological characterisation of graph links [18, Theorem 3.3] in this general setting.
@@ 5.3. Graph manifolds prime to $S^{1}\times S^{2}$
    Form the connected sum of the $M_{i}$ as in [4, Section 6].
@@ 5.4. Graph manifolds containing $(S^{1}\times S^{2})$-summands
represented by a graph link. On $N$, as before, we choose a Bott-integrable contact structure that is $H_{1}$-complete on the Seifert fibred pieces coming from the JSJ decompositions of the irreducible summands of $N$. On $k(S^{1}\times S^{2})$ we may choose, by [4, Theorem 1.9], a Bott-integrable contact structure whose Euler class has the Poincaré dual $b_{\eta}$. Write $\xi_{0}$ for the Bott-integrable contact structure on $M$ obtained via the connected sum.
@@ 6.3. The classification of Bott-integrable contact structures on $M(\psi)$
which is conjugate to $T^{-3}S$. Therefore, by Honda’s classification (up to isotopy) of tight contact structures on torus bundles, see specifically the table on page 90 of [9], all tight contact structures on $M(\psi)$ are universally tight, that is, their lift to $\mathbb{R}^{3}$ is tight. There is a unique tight structure for each value $2n\pi$, $n\in\mathbb{N}$, of its Giroux torsion [7, Definition 1.2], and two with zero Giroux torsion. For the relation between the Giroux torsion and the ‘twisting in the $S^{1}$-direction’ used by Honda, see [9, page 86].
Giroux’s classification [7] of tight contact structures on torus bundles is up to diffeomorphism, and according to [7, Theorem 1.3], there is a unique one for each value $2n\pi$ of the Giroux torsion, including the case $n=0$.
@@
On $\mathbb{R}^{3}$, the $1$-form $\alpha_{n}$, just like any $1$-form of Lutz type, defines the standard tight contact structure; see [6, Lemme 1] for the simple argument. For $n=0$, the coefficient function $h_{2}$ of $\mathrm{d}y$ is always positive, so the Giroux torsion of $\ker\alpha_{0}$ is zero. An increase of $n$ by $1$ introduces one additional full twist around the origin by the curve $t\mapsto(h_{1}(t),h_{2}(t))$ on the interval $t\in[0,1]$, which implies an additional $2\pi$ Giroux torsion.
As there is no $2$-torsion in $H^{2}(M(\psi))$, the overtwisted contact structures are classified by their Euler class and their $3$-dimensional homotopy invariant. As we have seen, the Bott-integrable ones must have trivial Euler class, so we can obtain all of them simply by taking the connected sum of $\bigl{(}M(\psi),\ker\alpha_{0}\bigr{)}$, say, with any of the overtwisted contact structures on $S^{3}$, which are Bott integrable by [4, Theorem 1.9]. ∎

### v2

@@ 4.1.2. Introducing Lutz components and critical orbits
This modified contact form still has the same Bott integral (a monotone function of $s$), and the homotopy class of the contact structure as a tangent $2$-plane field is unchanged by the full Lutz twist; see [5, Lemma 4.5.3].
In a second step, we modify the Bott integral by a finite number of ‘reverse Giroux eliminations’ in the $(s,\varphi)$- or the $(s,\theta)$-plane transverse to the Reeb flow in the direction of ±$\pm\partial_{\theta}$ or $\partial_{\varphi}$, respectively; see [6, Section 2.3] for the details. In this way, we can introduce any finite number of pairs of elliptic and hyperbolic critical orbits into the Reeb flow, lying in the isotopy class of ±$\pm S^{1}_{\theta}$ (the fibre class) or $S^{1}_{\varphi}$ (the class represented by $\gamma$).
@@ 4.2.2. Fibre connected sum
The general construction of fibre connected sums of contact manifolds along two diffeomorphic contact submanifolds with opposite normal bundles has been described in [5, Section 7.4].
In our situation, an ad hoc construction is much simpler and yields the Bott-integrability to boot. Indeed, the fibre connected sum along two elliptic critical Reeb orbits can be defined by removing standard neighbourhoods as in [6, Theorem 2.4] around the two orbits, and then identifying the toral boundaries of the complements by a diffeomorphism that sends the fibre circle on one side to the fibre circle on the other, and glues the meridional circles with a flip of the orientation. Notice that the outer normal on one complement glues with the inner normal on the other, so the gluing of the torus boundary should indeed be done with an orientation-reversing diffeomorphism.
Since the contact form near the boundary of a tubular neighbourhood of an elliptic orbit is of Lutz type, the Bott-integrable contact form on the fibre connected sum can now be constructed by an application of the sewing lemma [6, Lemma 3.4].
@@ 5.1. Graph manifolds and the JSJ decomposition
The summands $M_{1},\ldots,M_{n}$ in the prime decomposition of a graph manifold $M$ are also graph manifolds; see Theorems 6.3 and 7.1 in [21]. Conversely, the connected sum of graph manifolds is a graph manifold; see the remark on page 91 of [21] or [6, Lemma 3.1].
@@
If $M$ is an irreducible manifold, it can be cut along a disjoint collection of incompressible tori into pieces that are either atoroidal or Seifert fibred. Rather amazingly, the minimal (with respect to inclusion) collection of such tori is unique up to isotopy; see [10, Theorem 1.9]. This unique decomposition is known as the JSJ decomposition of $M$, named after Jaco–Shalen [12] and Johannson [14].
It turns out that amongst irreducible $3$-manifolds, graph manifolds are characterised as those whose JSJ decomposition only contains Seifert fibred pieces. One direction of this statement is obvious; the other is the content of [21, Theorem 6.3]. In fact, in the contemporary literature this characterisation is often taken as the definition of a graph manifold.
@@ 5.2.1. Graph manifolds prime to $S^{1}\times S^{2}$
As a crucial preliminary step for this theorem, Yano establishes the particular case that every element in the first homology group of a Seifert manifold can be represented by a graph link [23, Proposition 2.2]. Our argument in Section 4 gives an alternative proof of that part.
@@ 5.2.2. Graph manifolds containing $(S^{1}\times S^{2})$-summands
Here is Yano’s homological characterisation of graph links [23, Theorem 3.3] in this general setting.
@@ 5.3. Graph manifolds prime to $S^{1}\times S^{2}$
    Form the connected sum of the $M_{i}$ as in [6, Section 6].
@@ 5.4. Graph manifolds containing $(S^{1}\times S^{2})$-summands
represented by a graph link. On $N$, as before, we choose a Bott-integrable contact structure that is $H_{1}$-complete on the Seifert fibred pieces coming from the JSJ decompositions of the irreducible summands of $N$. On $k(S^{1}\times S^{2})$ we may choose, by [6, Theorem 1.9], a Bott-integrable contact structure whose Euler class has the Poincaré dual $b_{\eta}$. Write $\xi_{0}$ for the Bott-integrable contact structure on $M$ obtained via the connected sum.
@@ 6.3. The classification of Bott-integrable contact structures on $M(\psi)$
which is conjugate to $T^{-3}S$. Therefore, by Honda’s classification (up to isotopy) of tight contact structures on torus bundles, see specifically the table on page 90 of [11], all tight contact structures on $M(\psi)$ are universally tight, that is, their lift to $\mathbb{R}^{3}$ is tight. There is a unique tight structure for each value $2n\pi$, $n\in\mathbb{N}$, of its Giroux torsion [9, Definition 1.2], and two with zero Giroux torsion. For the relation between the Giroux torsion and the ‘twisting in the $S^{1}$-direction’ used by Honda, see [11, page 86].
Giroux’s classification [9] of tight contact structures on torus bundles is up to diffeomorphism, and according to [9, Theorem 1.3], there is a unique one for each value $2n\pi$ of the Giroux torsion, including the case $n=0$.
@@
On $\mathbb{R}^{3}$, the $1$-form $\alpha_{n}$, just like any $1$-form of Lutz type, defines the standard tight contact structure; see [8, Lemme 1] for the simple argument. For $n=0$, the coefficient function $h_{2}$ of $\mathrm{d}y$ is always positive, so the Giroux torsion of $\ker\alpha_{0}$ is zero. An increase of $n$ by $1$ introduces one additional full twist around the origin by the curve $t\mapsto(h_{1}(t),h_{2}(t))$ on the interval $t\in[0,1]$, which implies an additional $2\pi$ Giroux torsion.
As there is no $2$-torsion in $H^{2}(M(\psi))$, the overtwisted contact structures are classified by their Euler class and their $3$-dimensional homotopy invariant. As we have seen, the Bott-integrable ones must have trivial Euler class, so we can obtain all of them simply by taking the connected sum of $\bigl{(}M(\psi),\ker\alpha_{0}\bigr{)}$, say, with any of the overtwisted contact structures on $S^{3}$, which are Bott integrable by [6, Theorem 1.9]. ∎